Grade 12th passMechanics

A body moves along a straight line with acceleration 3m/s^2 for 2s and then with acceleration 4m/s^2for 3s.what is the average acceleration?

Shreya Ghosh
8 Years agoGrade 12th pass

1 Answer

Khimraj
8 Years ago
Velocity after 2s 
v1 = 3*2 = 6m/s
velocity after 5s
v2 = 6 + 4*3 = 18m/s
Average acceleration = (change in velocity)/(time taken) = (v2 -0)/(5-0) = 18/5 = 3.6m/s2.
Hope it clears. If you like answer then please approve it.